# Detecting Early and Implicit Suicidal Ideation via Longitudinal and Information Environment Signals on Social Media

### v1

@@ 4.2 Neighboring user detection
$C_{\text{in-degree}}$ and $C_{\text{out-degree}}$ capture normalized connectivity, $C_{\text{betweenness}}$ measures shortest-path centrality, $C_{\text{closeness}}$ denotes proximity, $C_{\text{eigenvector}}$ reflects influence via important neighbors, and $C_{\text{PageRank}}$ estimates probabilistic importance. The aggregated NeighborScore identifies peers with strong direct ties and broader network influence around $U$. As some centralities are correlated, $S(n)$ approximates overall neighbor prominence rather than precise influence. Performance variation by neighbor depth and count is shown in Appendix Table A9.
@@ 4.4 Implicit SI signal classification with modeling
For our study, we leveraged Microsoft’s Deberta-v3-large model He et al. (2021)—a 418M parameter transformer with 24 layers that processes sequences up to 512 tokens. We fine-tuned this model using CLS for global representation and SEP for segment boundaries on an Nvidia A100 GPU. We framed our problem of detecting implicit SI as a binary classification task—labeling each input $x_{i}$ as $y_{i}\!\in\!\{0,1\}$ for absence or presence of risk. We tokenized text into subword embeddings with attention masks, padded or truncated to 512 tokens, and encoded them to obtain a pooled [CLS] vector (1024-dim). A linear layer then mapped this vector to logits, followed by softmax for class probabilities. We optimized the model with cross-entropy loss using AdamW (learning rate $2\times 10^{-5}$, weight decay $0.01$) for 20 epochs with batch size 8, and evaluated after each epoch on a held-out validation set using accuracy, precision, recall, and F1-score. Results from fine-tuning alternative base models are provided in Appendix Table A2.
@@ 5.4 Robustness Tests:
To ensure that our findings are not artifacts of specific modeling or sampling choices, we conducted additional robustness tests by varying the type of neighbors included—for instance, incorporating non-top neighbors into our models. For this purpose, we built additional models $M_{5}$ with the lowest-ranked neighbors and $M_{6}$ with non-neighbors—a random selection of other users—of a target user. Table 2 compares the performance—we note both $M_{4}$ and $M_{5}$ perform significantly poorly. Therefore, the performance boost by including the top neighbor posts is not by chance, and rather an important element in detecting implicit SI Cero et al. (2024). Moreover, Table 3 demonstrates that the model retains strong predictive performance even after filtering out all top neighbors who have posted in r/SuicideWatch. This finding indicates that successful detection does not rely solely on neighbors’ direct SI-related language, but rather on broader contextual cues captured from high-quality interactions. Appendix Table A4 shows linguistic analyses aligned with Table 2, Table A3 highlights “Affect” differences linked to Table 3, and additional model comparisons with LIWC results are in Appendix Table A1 (Tables A5A8).
@@ 10 AI Involvement Disclosure
Appendix A Appendix
Table A1: Model Performance on Remaining Data Combinations (epochs=20).
Model Data Used Accuracy F1 Precision Recall
$M_{8}$ Top neighbor’s posts only 0.83 0.82 0.92 0.74
$M_{9}$ Target user posts + comments + top neighbor posts 0.91 0.92 0.91 0.92
$M_{10}$ Target user posts + comments + others’ comments + top neighbor posts 0.71 0.64 0.75 0.56
$M_{11}$ Filtered neighbor posts only (no post/comments in r/SuicideWatch) 0.78 0.77 0.73 0.82
Table A2: Performance comparison of transformer models on the classification task.
Model Accuracy F1 Precision Recall
microsoft/deberta-v3-large 0.955 0.946 0.945 0.945
roberta-large 0.944 0.931 0.944 0.919
bert-base-uncased 0.933 0.921 0.897 0.946
google/electra-large-discriminator 0.944 0.935 0.900 0.973
Table A3: Comparison of LIWC affect categories between $M_{4}$ and $M_{7}$ showing how the LIWC Affect category varies when considering all top neighbors vs. filtered top neighbors. Columns show mean proportions for users showing suicidal ideation (SI) and not, their difference, and the Kruskal–Wallis H-statistic. Significance codes: ***pp<0.001, **pp<0.01, *pp<0.05.
Category $M_{4}$ $M_{7}$
Case Control Diff. H Case Control Diff. H
posemo 0.048 0.053 -0.005 23.568*** 0.050 0.053 -0.003 17.482***
negemo 0.029 0.021 0.007 169.871*** 0.032 0.025 0.007 72.531***
anx 0.003 0.002 0.000 62.663*** 0.003 0.002 0.001 63.084***
anger 0.011 0.006 0.005 130.857*** 0.010 0.006 0.004 51.436***
sad 0.009 0.005 0.004 181.551*** 0.008 0.007 0.001 71.131***
Table A4: Comparison of LIWC affect categories across $M_{4}$, $M_{5}$, and $M_{6}$ showing how the Affect LIWC category varies across various neighbors (top, worst, random). Columns show mean proportions for users showing suicidal ideation (SI) and not, their difference, and the Kruskal–Wallis H-statistic. Significance codes: ***pp<0.001, **pp<0.01, *pp<0.05.
Category $M_{4}$ $M_{5}$ $M_{6}$
Case Control Diff. H Case Control Diff. H Case Control Diff. H
posemo 0.048 0.053 -0.005 23.568*** 0.047 0.049 -0.003 38.611*** 0.048 0.047 0.001 42.425***
negemo 0.029 0.021 0.007 169.871*** 0.029 0.025 0.004 74.947*** 0.030 0.024 0.006 59.078***
anx 0.003 0.002 0.000 62.663*** 0.004 0.002 0.002 53.518*** 0.003 0.002 0.001 125.397***
anger 0.011 0.006 0.005 130.857*** 0.009 0.006 0.003 40.154*** 0.009 0.006 0.002 138.704***
sad 0.009 0.005 0.004 181.551*** 0.007 0.007 0.000 71.131*** 0.007 0.008 -0.001 36.640***
Table A5: Comparison of LIWC affect categories for $M_{8}$. Columns show mean proportions for suicidal and non-suicidal users, their difference, and the Kruskal–Wallis $H$ statistic. Significance codes: $^{***}p<0.001$, $^{**}p<0.01$, $^{*}p<0.05$.
Category $M_{8}$
Case Control Diff H
posemo 0.053 0.053 0.000 0.409
negemo 0.026 0.025 0.001 4.805*
anx 0.001 0.002 -0.001 3.323
anger 0.009 0.006 0.003 9.523**
sad 0.005 0.007 -0.001 3.638
Table A6: Comparison of LIWC affect categories for $M_{9}$. Columns show mean proportions for suicidal and non-suicidal users, their difference, and the Kruskal–Wallis $H$ statistic. Significance codes: $^{***}p<0.001$, $^{**}p<0.01$, $^{*}p<0.05$.
Category $M_{9}$
Case Control Diff H
posemo 0.050 0.053 -0.003 17.482***
negemo 0.032 0.025 0.007 72.531***
anx 0.003 0.002 0.001 63.084***
anger 0.010 0.006 0.004 51.436***
sad 0.008 0.007 0.001 71.131***
Table A7: Comparison of LIWC affect categories for $M_{10}$. Columns show mean proportions for suicidal and non-suicidal users, their difference, and the Kruskal–Wallis $H$ statistic. Significance codes: $^{***}p<0.001$, $^{**}p<0.01$, $^{*}p<0.05$.
Category $M_{10}$
Case Control Diff H
posemo 0.070 0.067 0.003 349.268***
negemo 0.031 0.028 0.003 538.074***
anx 0.003 0.002 0.000 403.691***
anger 0.010 0.009 0.001 281.206***
sad 0.007 0.006 0.001 361.992***
Table A8: Comparison of LIWC affect categories for $M_{11}$. Columns show mean proportions for suicidal and non-suicidal users, their difference, and the Kruskal–Wallis $H$ statistic. Significance codes: $^{***}p<0.001$, $^{**}p<0.01$, $^{*}p<0.05$.
Category $M_{11}$
Case Control Diff H
posemo 0.053 0.053 0.000 0.409
negemo 0.026 0.025 0.001 4.805*
anx 0.001 0.002 -0.001 3.323
anger 0.009 0.006 0.003 9.523**
sad 0.005 0.007 -0.001 3.638
Table A9: Performance metrics for different depth and neighbor configurations. Depth refers to the number of recursive depth considered, and Neighbors refers to the number of top neighbor posts included.
Depth Neighbors Accuracy F1 Precision Recall
1 1 0.98 0.97 0.96 0.98
1 2 0.97 0.96 0.95 0.97
1 3 0.96 0.95 0.95 0.96
1 4 0.96 0.95 0.94 0.96
1 5 0.95 0.94 0.94 0.95
2 1 0.97 0.96 0.95 0.97
2 2 0.96 0.95 0.94 0.96
2 3 0.95 0.94 0.93 0.95
2 4 0.95 0.94 0.93 0.95
2 5 0.94 0.93 0.92 0.94
3 1 0.96 0.95 0.94 0.96
3 2 0.95 0.94 0.93 0.95
3 3 0.94 0.93 0.92 0.94
3 4 0.94 0.93 0.92 0.94
3 5 0.95 0.94 0.93 0.95
4 1 0.95 0.94 0.93 0.95
4 2 0.94 0.93 0.92 0.94
4 3 0.93 0.92 0.91 0.93
4 4 0.93 0.92 0.91 0.93
4 5 0.92 0.91 0.90 0.92
5 1 0.94 0.93 0.92 0.94
5 2 0.93 0.92 0.91 0.93
5 3 0.92 0.91 0.90 0.92
5 4 0.92 0.91 0.90 0.92
5 5 0.91 0.90 0.89 0.91

### v2

@@ 4.2 Neighboring user detection
$C_{\text{in-degree}}$ and $C_{\text{out-degree}}$ capture normalized connectivity, $C_{\text{betweenness}}$ measures shortest-path centrality, $C_{\text{closeness}}$ denotes proximity, $C_{\text{eigenvector}}$ reflects influence via important neighbors, and $C_{\text{PageRank}}$ estimates probabilistic importance. The aggregated NeighborScore identifies peers with strong direct ties and broader network influence around $U$. As some centralities are correlated, $S(n)$ approximates overall neighbor prominence rather than precise influence.
@@ 4.4 Implicit SI signal classification with modeling
For our study, we leveraged Microsoft’s Deberta-v3-large model He et al. (2021)—a 418M parameter transformer with 24 layers that processes sequences up to 512 tokens. We fine-tuned this model using CLS for global representation and SEP for segment boundaries on an Nvidia A100 GPU. We framed our problem of detecting implicit SI as a binary classification task—labeling each input $x_{i}$ as $y_{i}\!\in\!\{0,1\}$ for absence or presence of risk. We tokenized text into subword embeddings with attention masks, padded or truncated to 512 tokens, and encoded them to obtain a pooled [CLS] vector (1024-dim). A linear layer then mapped this vector to logits, followed by softmax for class probabilities. We optimized the model with cross-entropy loss using AdamW (learning rate $2\times 10^{-5}$, weight decay $0.01$) for 20 epochs with batch size 8, and evaluated after each epoch on a held-out validation set using accuracy, precision, recall, and F1-score.
@@ 5.4 Robustness Tests:
To ensure that our findings are not artifacts of specific modeling or sampling choices, we conducted additional robustness tests by varying the type of neighbors included—for instance, incorporating non-top neighbors into our models. For this purpose, we built additional models $M_{5}$ with the lowest-ranked neighbors and $M_{6}$ with non-neighbors—a random selection of other users—of a target user. Table 2 compares the performance—we note both $M_{4}$ and $M_{5}$ perform significantly poorly. Therefore, the performance boost by including the top neighbor posts is not by chance, and rather an important element in detecting implicit SI Cero et al. (2024). Moreover, Table 3 demonstrates that the model retains strong predictive performance even after filtering out all top neighbors who have posted in r/SuicideWatch. This finding indicates that successful detection does not rely solely on neighbors’ direct SI-related language, but rather on broader contextual cues captured from high-quality interactions.